on a recent math test , the mean score was 75 and the standard deviation was 5. mike got 85. what percentile does he fall in?


1
Expert's answer
2022-03-29T05:30:43-0400

"Z=\frac{x-\mu}{\sigma}=\frac{85-75}{5}=2"

P(x<85)=0.97725